OpenStudy (anonymous):

Evaluate 5m – 2n for m = –3 and n = –2. (1 point) A.–19 B.–11 C.19 D.11

OpenStudy (a_clan):

Substitute the values of m and n in the given expression. 5(-3) - 2(-2) First you multiply brackets =(-15) - (-4) =(-15) + 4 = -15 + 4 =-11 (ans)